Exercise 21.57 ia Exercise 21.57 Point charges h =-4.20 nC and g2 4.20 nC are separated by a distance of 4.00 mm, forming an electric dipole
Answer #1

a) Elctric dipole moment (p) = q x d

p = 4.20 x 10-9 C x 4 x 10-3 m = 1.68 x 10-11 C m

b)   q1 to q2

c) Torque = p x E = p E sin(36.6o)

== 7.4 x 10 -9 = 1.68 x 10 -11 x E x 0.596

E = 7.39 x 10 2 N/C
Answer #2

SOLUTION :


a.


q = magnitude of the couple formed = 4.20 nC 


Electric Di-pole moment  (magnitude) of given charges of opposite but equal magnitude :

=  magnitude of  charge in C *  distance in m

= 4.20 * 10^(- 9) * 4 * 10^(-3)

= 16.80 * 10^(-12)

= 1.68 * 10^(-11)     (C-m) (ANSWER).


b.


Direction of the electric dipole moment is from negative charge to positive charge.


So, this direction is from q1 to q2. (ANSWER).


c.


Torque magnitude 

= Polar moment magnitude * Electric field E magnitude  * sin (angle between them)

=> 7.40 * 10(-9)  = 1.68*10(-11) * E (magnitude) * sin(36.6)

=> E (magnitude) = 7.40 * 10^(-9) / (1.68 * 10^(-11) * sin(36.6))  

=> E (magnitude) = 738.78 N/C (ANSWER). 

answered by: Tulsiram Garg
Answer #3

SOLUTION :


a.


q = magnitude of the couple formed = 4.20 nC 


Electric Di-pole moment  (magnitude) of given charges of opposite but equal magnitude :

=  magnitude of  charge in C *  distance in m

= 4.20 * 10^(- 9) * 4 * 10^(-3)

= 16.80 * 10^(-12)

= 1.68 * 10^(-11)     (C-m) (ANSWER).


b.


Direction of the electric dipole moment is from negative charge to positive charge.


So, this direction is from q1 to q2. (ANSWER).


c.


Torque magnitude 

= Polar moment magnitude * Electric field E magnitude  * sin (angle between them)

=> 7.40 * 10(-9)  = 1.68*10(-11) * E (magnitude) * sin(36.6)

=> E (magnitude) = 7.40 * 10^(-9) / (1.68 * 10^(-11) * sin(36.6))  

=> E (magnitude) = 738.78 N/C (ANSWER). 

answered by: Tulsiram Garg